# A Scalable and Parallelizable Digital Twin Framework for Sustainable Sim2Real Transition of Multi-Agent Reinforcement Learning Systems

### v4

@@ Abstract
Multi-agent reinforcement learning (MARL) systems usually require significantly long training times due to their inherent complexity. Furthermore, deploying them in the real world demands a feature-rich environment along with multiple embodied agents, which may not be feasible due to budget or space limitations, not to mention energy consumption and safety issues. This work tries to address these pain points by presenting a sustainable digital twin framework capable of accelerating MARL training by selectively scaling parallelized workloads on-demand, and transferring the trained policies from simulation to reality using minimal hardware resources. The applicability of the proposed digital twin framework is highlighted through two representative use cases, which cover cooperative as well as competitive classes of MARL problems. We study the effect of agent and environment parallelization on training time and that of systematic domain randomization on zero-shot sim2real transfer across both the case studies. Results indicate up to 76.3% reduction in training time with the proposed parallelization scheme and as low as 2.9% sim2real gap using the suggested deployment method.
@@ I Introduction
In the rapidly evolving landscape of connected and autonomous vehicles (CAVs), the pursuit of intelligent and adaptive driving systems has emerged as a formidable challenge. In such a milieu, multi-agent learning stands out as a promising avenue for developing autonomous vehicles capable of navigating complex and dynamic environments while considering the nature of interactions with their peers. Particularly, multi-agent reinforcement learning (MARL) offers the tantalizing potential of learning through self-exploration, which can potentially capture intricate cooperative/competitive multi-agent interactions.
@@
Figure 1: Resource-aware digital twin framework for sim2real transfer of multi-agent deep reinforcement learning systems: (1) observe, (2) decide, (3) act, (4) estimate, and (5) update.
@@
Cooperative MARL [1, 2, 3, 4, 5, 6, 7] fosters an environment where autonomous vehicles collaborate and share information to accomplish collective objectives such as optimizing traffic flow, enhancing safety, and efficiently navigating road networks. It mirrors traffic situations where vehicles must work together, such as intersection management or platooning scenarios. Challenges in cooperative MARL include coordinating vehicle actions to minimize congestion, maintaining safety margins, and ensuring smooth interactions with peers.
On the other hand, competitive MARL [8, 9, 10, 11, 12] introduces elements of privacy and rivalry in autonomous driving, simulating scenarios such as overtaking, merging in congested traffic, or racing. In this paradigm, agents strive to outperform their counterparts, prioritizing individual success over coordination. Challenges in competitive MARL encompass strategic decision-making, opponent modeling, and adapting to aggressive driving behaviors while preserving safety.
Irrespective of the problem formulation, one of the key challenges in training MARL policies is the sample efficiency of the environment, which translates to longer training times. This is often addressed by (a) training in low-fidelity simulations that can run faster than real-time [2, 3, 4, 5, 6], (b) parallelizing simulations to accelerate data collection [9, 10], or (c) both [12]. Adopting the first approach usually leads to a heightened sim2real gap, as marked in the literature through simulation-only deployments or explicit remarks for real-world experiments [12]. Adopting the second approach, on the other hand, usually requires extensive computational resources [9, 10, 13], which may not be sustainable. Additionally, none of the prior works addresses the challenge of selectively isolating collision, interaction, and perception between parallelized replicas of MARL systems.
Another important challenge arises during the real-world deployment of MARL policies. While generic RL has studied the sim2real transfer of trained policies using techniques such as domain adaptation [14, 15, 16], identification [17, 18, 19], or augmentation [20, 21, 22, 19], the sim2real transfer of MARL systems has been typically under-explored. While recent works such as [7] and [12] filled this gap, they simplified problem formulation by adopting extensive observation spaces with ground truth information about the environment and employed benchmark equipment such as mocap for real-time feedback during sim2real transfer. Additionally, these works used multiple physical vehicles, albeit scaled, within a synthetically constructed physical test environment for real-world deployments, which may not be sustainable.
@@
  • MARL Analysis: The proposed MARL formulations are benchmarked against a decentralized reactive planning algorithm to assess their efficacy. We numerically evaluate the sim2real gap to analyze the effect of systematic domain randomization introduced in this work.
II Formulation
@@ II-A Cooperative Multi-Agent Scenario
This section formulates a decentralized 4-agent collaborative scenario, wherein each agent’s objective was to traverse a 2+2 lane, 4-way intersection without colliding or overstepping lane bounds. Each agent perceived its intrinsic states and received limited information from its peers (via V2V communication); no external sensing modalities were employed. Each agent was reset independently, resulting in highly stochastic initial conditions. The exact structure/map of the environment was not known to any agent. Consequently, this problem was framed as a partially observable Markov decision process (POMDP), which captured hidden state information through limited observations.
@@ II-A2 Action Space
The Ackermann-steered vehicles were controlled using discrete throttle and steering commands: $a_{t}^{i}=\left[\tau_{t}^{i},\delta_{t}^{i}\right]\in\mathbb{R}^{2}$. Here, the throttle command $\tau_{t}\in\left\{0.5,1.0\right\}$ allowed the agents to slow down if necessary, and steering command $\delta_{t}\in\left\{-1,0,1\right\}$ allowed the agents to steer left, straight, or right for collision avoidance.
@@ II-B Competitive Multi-Agent Scenario
This section formulates a 2-agent autonomous racing scenario, wherein each agent’s objective was to minimize its lap time without colliding with the track or its opponent. Each agent collected its observations; no information was shared among the agents. The exact map of the environment was not known to any agent. Consequently, this problem was also framed as a POMDP. However, contrary to the cooperative MARL, this problem adopted a hybrid imitation-reinforcement learning architecture to guide the agents’ exploration, thereby reducing training time. To this end, we recorded 5 laps worth of single-agent demonstrations for each agent by manually driving it in sub-optimal trajectories.
@@ II-B2 Action Space
The action space to control Ackermann-steered vehicles was $a_{t}^{i}=\left[\tau_{t}^{i},\delta_{t}^{i}\right]\in\mathbb{R}^{2}$. Here, $\tau_{t}^{i}\in\left\{0.1,0.5,1.0\right\}$ represents the discrete throttle commands for torque-limited (85.6 N-m) actuation, and $\delta_{t}^{i}\in\left\{-1,0,1\right\}$ represents the left, straight, and right steering commands.
@@ III Methodology
In this work, we adopted and adapted the AutoDRIVE Ecosystem [26, 27] to model, simulate, train, and deploy two MARL case studies. Particularly, we created analytical models of Nigel [28] and F1TENTH [29] complemented with data-driven system identification and calibration. Further details about vehicle, sensor, and environment modeling within AutoDRIVE Simulator [30, 31] are available in [32]. That being said, this section describes certain novel enhancements to AutoDRIVE Ecosystem introduced in this work.
@@ III-A Simulation Parallelization
From a computing perspective, AutoDRIVE Simulator was developed modularly using object-oriented programming (OOP) constructs. We built on this fact to implement a selectively scalable agent/environment parallelization architecture. The simulator was configured to take advantage of CPU multi-threading as well as GPU instancing (only if available) to efficiently parallelize various simulation objects and processes while maintaining cross-platform support. Following is an overview of the simulation parallelization schemes:
@@ III-B Learning Architecture
We leverage the proximal policy optimization (PPO) algorithm [33] to train both the multi-agent systems described in this work; justification follows. PPO is an on-policy method, which is empirically equally competitive to its off-policy counterparts [34]. Additionally, PPO promotes stable and efficient learning by imposing 2 complementary constraints: (a) a clipped surrogate objective to control each action probability update, and (b) a KL divergence early stopping criteria to limit overall policy change.
@@
Table I hosts the detailed training configurations adopted for the cooperative as well as competitive MARL scenarios. The parameter values mentioned were arrived at by analyzing the agent(s)’ behaviors to satisfy the intended objectives qualitatively, while also ensuring a stable learning process. All the trainings were carried out on a single laptop PC having 12th Gen Intel Core i9-12900H 2.50 GHz CPU, NVIDIA GeForce RTX 3080 Ti GPU, and 32.0 GB RAM.
@@ III-C Domain Randomization
III-D Sustainable Sim2Real Transfer
Our work proposes a sustainable method of transferring the trained MARL policies from simulation to reality. The term “sustainability”, here specifically alludes to resource-awareness, where we aim to minimize the physical resources in deploying and validating MARL systems.
To this end, we present a bi-directional digital twinning framework, which establishes a real-time synchronization between the physical and virtual realms. Fig. 1 (captured at 1 Hz) depicts the sim2real transfer of MARL policies discussed in this work using the proposed digital twinning framework. Here, we deploy a single physical agent in an open space and connect it with its digital twin, which operates in a virtual environment along with other virtual agents to generate actions in the digital space. These action sequences are then relayed back to the physical agent to be executed in the real world. This way, we can exploit the real-world characteristics of vehicle dynamics and tire-road interactions while being resource-altruistic by augmenting environmental element(s) and peer agent(s) in the digital space.
@@ IV-A1 Training and Simulation Parallelization
Figure 3: Training results for cooperative MARL: (a) denotes cumulative reward, (b) denotes episode length, and (c) denotes policy entropy w.r.t. training steps.
@@ IV-A2 Deployment and Sim2Real Transfer
The trained policies were first deployed and verified in simulation, where we observed interesting emergent behaviors among the agents. The agents strategically slowed down and/or steered away from each other to avoid collision.
@@
Finally, it is worth mentioning that the closeness between sim and real metrics effectively measures the sim2real gap, which was least for LDR-MARL (4.12%) followed by NDR-MARL (9.38%), FGM (16.01%), and HDR-MARL (33.85%) respectively.
@@ IV-B Competitive Multi-Agent Scenario
Figure 5: Training results for competitive MARL: (a) denotes BC loss, (b) denotes GAIL reward, (c) denotes curiosity reward, (d) denotes extrinsic reward, (e) denotes episode length, and (f) denotes policy entropy w.r.t. training steps.
@@ IV-B2 Deployment and Sim2Real Transfer
The trained policies were first deployed and verified in simulation, where we observed interesting adversarial behaviors among the agents. One such behavior was a block-block-overtake sequence, and another was a let-pass-and-overtake sequence. These behaviors conveyed that the agents were explicitly competing, while implicitly coordinating to avoid collisions.
@@
Finally, the sim2real gap, which was measured by averaging the difference between sim and real performance across all KPIs, was least for LDR-MARL (2.88%) followed by NDR-MARL (6.88%), HDR-MARL (8.98%), and FGM (13.48%) respectively.
@@ V Conclusion
This work identified two pain points in training and deploying MARL systems, and attempted to address them by proposing a scalable and parallelizable digital twin framework. Two representative case studies were formulated to support the claims: a 4-agent collaborative intersection traversal problem and a 2-agent adversarial head-to-head racing problem. The two problems were deliberately formulated with distinct observation spaces and reward functions, but more importantly, also the learning architecture (vanilla MARL vs. demonstration-guided MARL). We analyzed the training metrics in each case and also noted the non-linear effect of agent/environment parallelization on the training time, with a hardware/software-specific point of diminishing return. Finally, we presented a sustainable sim2real transfer of the trained policies using a single physical vehicle, which was immersed within the proposed digital twin framework to interact with its virtual peers in a virtual environment.
One of the bottlenecks of the proposed framework is its dependence on the state estimation and communication protocols used to update the digital twins. High estimation errors or communication delays can affect the MARL policies to not yield the right actions at the right time. We intend to investigate this in the future. Other avenues of research include the formulation of physics-guided MARL problems and scaling the deployments in terms of the number and size of the agents.

### v5

# Mixed-Reality Digital Twins: Leveraging the Physical and Virtual Worlds for Hybrid Sim2Real Transition of Multi-Agent Reinforcement Learning Policies

@@ Abstract
Multi-agent reinforcement learning (MARL) for cyber-physical vehicle systems usually requires a significantly long training time due to their inherent complexity. Furthermore, deploying the trained policies in the real world demands a feature-rich environment along with multiple physical embodied agents, which may not be feasible due to monetary, physical, energy, or safety constraints. This work seeks to address these pain points by presenting a mixed-reality digital twin framework capable of: (i) selectively scaling parallelized workloads on-demand, and (ii) evaluating the trained policies across simulation-to-reality (sim2real) experiments. The viability and performance of the proposed framework are highlighted through two representative use cases, which cover cooperative as well as competitive classes of MARL problems. We study the effect of: (i) agent and environment parallelization on training time, and (ii) systematic domain randomization on zero-shot sim2real transfer across both case studies. Results indicate up to 76.3% reduction in training time with the proposed parallelization scheme and sim2real gap as low as 2.9% using the proposed deployment method.
Index Terms:
Multi-Agent Systems, Autonomous Vehicles, Reinforcement Learning, Digital Twins, Real2Sim, Sim2Real
@@ I Introduction
Connected autonomous vehicles (CAVs) are exemplars of cyber-physical systems (CPS) operating within an environment with other agents. The development and deployment of such systems present a formidable challenge due to the complexity-growth of multi-agent interactions. In such a milieu, multi-agent learning stands out as a promising avenue for developing autonomous vehicles capable of navigating complex and dynamic environments while considering the nature of interactions with their peers. Particularly, multi-agent reinforcement learning (MARL) offers the tantalizing potential of learning through self-exploration, which can potentially capture intricate cooperative/competitive multi-agent interactions.
@@
Figure 1: Mixed-reality digital twin framework for hybrid sim2real transfer of MARL systems: (1) observe, (2) decide, (3) act, (4) estimate, and (5) update.
@@
Cooperative MARL [1, 2, 3, 4] fosters an environment where autonomous vehicles collaborate and share information to accomplish collective objectives such as optimizing traffic flow, enhancing safety, and efficiently navigating road networks. It mirrors traffic situations where vehicles must work together, such as intersection management or platooning scenarios. Challenges in cooperative MARL include coordinating vehicle actions to minimize congestion, maintaining safety margins, and ensuring smooth interactions with peers.
On the other hand, competitive MARL [5, 6, 7, 8] introduces elements of privacy and rivalry in autonomous driving, simulating scenarios such as overtaking, merging in congested traffic, or racing. In this paradigm, agents strive to outperform their counterparts, prioritizing individual success over coordination. Challenges in competitive MARL encompass strategic decision-making, opponent modeling, and adapting to aggressive driving behaviors while preserving safety.
Irrespective of the problem formulation, one of the key challenges in training MARL policies is the sample efficiency of the environment, which translates to longer training times. This is often addressed by (a) training in low-fidelity simulations that can run faster than real-time [1, 2, 3], (b) parallelizing simulations to accelerate data collection [7, 8], or (c) both [6]. Adopting the first approach usually leads to a heightened sim2real gap, as marked in the literature through simulation-only deployments or explicit remarks for real-world experiments [6]. Adopting the second approach, on the other hand, usually requires extensive computational resources [7, 8, 9], which may not be sustainable. Additionally, none of the prior works addresses the challenge of selectively isolating collision, interaction, and perception between parallelized replicas of multi-agent systems.
Another important challenge arises during the real-world deployment of MARL policies. While generic RL has studied the sim2real transfer of trained policies using techniques such as domain adaptation [10], identification [11], or augmentation [12], the sim2real transfer of MARL systems has been typically under-explored. While recent works such as [4] and [6] have tried to fill this gap, they simplified the formulation by adopting extensive observation spaces with ground truth information about the environment and employed benchmark equipment such as mocap for real-time feedback during sim2real transfer. Additionally, these works used multiple physical vehicles, albeit scaled, within a synthetically constructed physical test environment for real-world deployments, which may not always be feasible due to monetary, spatial, energy, or safety constraints.
@@
  • MARL Analysis: The proposed MARL case studies are benchmarked against a decentralized reactive planning algorithm to assess their efficacy. We numerically evaluate the sim2real gap to analyze the effect of systematic domain randomization introduced in this work.
The remainder of this paper is organized as follows: Section II elucidates the two MARL case studies, including their mathematical formulation. Section III describes the workflow adopted to train and deploy the MARL policies, including their sim2real transfer. Section IV analyzes the MARL training and deployment results for both case studies. Finally, Section V provides concluding remarks and future research directions.
II Case Studies
@@ II-A Cooperative Multi-Agent Scenario
We formulated a decentralized 4-agent collaborative scenario, wherein each agent’s objective was to traverse a 2+2 lane, 4-way intersection without colliding or overstepping lane bounds. This scenario is representative of a standard uncontrolled traffic intersection. Here, each agent perceived its intrinsic states and received limited information from its peers (via V2V communication); no external sensing modalities were employed. Each agent was reset independently, resulting in highly stochastic initial conditions. The exact structure/map of the environment was not known to any agent. Consequently, this problem was framed as a partially observable Markov decision process (POMDP), which captured hidden state information through limited observations.
@@ II-A2 Action Space
The Ackermann-steered vehicles were controlled using throttle and steering commands: $a_{t}^{i}=\left[\tau_{t}^{i},\delta_{t}^{i}\right]\in\mathbb{R}^{2}$. From a collision avoidance perspective, the throttle command allowed the agents to speed up or slow down, and the steering command allowed the agents to dodge their peers.
@@ II-B Competitive Multi-Agent Scenario
We formulated a 2-agent autonomous racing scenario, wherein each agent’s objective was to minimize its lap time without colliding with the track or its opponent. This scenario is representative of a standard F1TENTH (a.k.a. RoboRacer) autonomous racing competition [13]. Here, each agent collected its observations; no information was shared among the agents. The exact map of the environment was not known to any agent. Consequently, this problem was also framed as a POMDP. However, contrary to the cooperative MARL, this problem adopted a hybrid imitation-reinforcement learning architecture to guide the agents’ exploration, thereby reducing training time. To this end, we recorded 5 laps worth of single-agent demonstrations for each agent by manually driving it in sub-optimal trajectories.
@@ II-B2 Action Space
The action space to control Ackermann-steered vehicles was $a_{t}^{i}=\left[\tau_{t}^{i},\delta_{t}^{i}\right]\in\mathbb{R}^{2}$. Here, the throttle command allowed the agents to optimize their speed profile, and the steering command allowed the agents to optimize their race line, overtake their peers, and avoid collisions.
@@ III Methodology
In this work, we adopted and adapted the AutoDRIVE Ecosystem [17] to model, simulate, train, and deploy two MARL case studies. This choice was driven based on the comparative analysis presented in [17], which satisfied all the requirements of this study. From a digital twinning perspective, data-driven system identification and calibration were used to customize models of Nigel [18] and F1TENTH [19] vehicles from real-world data to ensure reliable simulation.
@@ III-A Simulation Parallelization
We leveraged the open-source nature of AutoDRIVE Simulator to implement a selectively scalable agent/environment parallelization framework. The simulator was configured to take advantage of CPU multi-threading as well as GPU instancing (only if available) to efficiently parallelize various simulation objects and processes while maintaining cross-platform support. Following is an overview of the simulation parallelization schemes:
@@ III-B Learning Architecture
We leverage the proximal policy optimization (PPO) algorithm [20] for MARL training; justification follows. PPO is an on-policy method, which is empirically equally effective as its off-policy counterparts [21]. Moreover, PPO promotes stable and efficient learning by imposing 2 complementary constraints: (a) a clipped surrogate objective to control each action probability update, and (b) a KL divergence early stopping criteria to limit overall policy change.
@@
Table I hosts the detailed training configurations adopted for the cooperative as well as competitive MARL scenarios. The noted parameter values were arrived at by analyzing the agent(s)’ behaviors to satisfy the intended objectives qualitatively, while also ensuring a stable learning process. MARL training was carried out on a single laptop PC with 12th Gen Intel Core i9-12900H 2.50 GHz CPU, NVIDIA GeForce RTX 3080 Ti GPU, and 32.0 GB RAM.
@@ III-C Domain Randomization
III-D Hybrid Sim2Real Transfer
We propose a hybrid method for transferring the trained MARL policies from simulation to reality. The term “hybrid” specifically alludes to a mixed-reality digital twin framework, which establishes a real-time bi-directional synchronization between the physical and virtual worlds. The intention is to minimize the number of physical agent(s) and environmental element(s) while deploying and validating MARL systems in the real world. Fig. 1 (captured at 1 Hz) depicts the sim2real transfer of MARL policies discussed in this work using the proposed framework. Here, we deploy a single physical agent in an open space and connect it with its digital twin. The “ego” digital twin operates in a virtual environment with virtual peers, collects observations, and uses the trained policy to plan actions in the digital space. The planned action sequences are relayed back to the physical twin to be executed in the real world, which updates its state in reality. Finally, the ego digital twin is updated based on real-time state estimates of its physical twin to close the loop. This process is repeated recursively until the experiment is completed. This way, we can exploit the real-world characteristics of vehicle dynamics and tire-road interactions while being resource-altruistic by augmenting environmental element(s) and peer agent(s) in the digital space. This also alleviates the safety concern of the experimental vehicles colliding with each other or the environmental element(s).
@@ IV-A1 Training and Simulation Parallelization
Figure 3: Training results for cooperative MARL: (a) cumulative reward, (b) episode length, and (c) policy entropy w.r.t. training steps.
Figure 4: Deployment results for cooperative MARL: (a) A1 and A4 avoid collision, (b) A1 finds a gap between A2 and A3 to reach its goal, and (c) A2 and A3 avoid collision, A4 approaches its goal, and A1 is re-spawned.
@@
Figure 6: Training results for competitive MARL: (a) BC loss, (b) GAIL reward, (c) curiosity reward, (d) extrinsic reward, (e) episode length, and (f) policy entropy w.r.t. training steps.
Figure 7: Deployment results for competitive MARL: (a)-(c) denote three frozen snapshots of a block-block-overtake sequence, and (d)-(f) denote three frozen snapshots of a let-pass-and-overtake sequence.
@@ IV-A2 Deployment and Sim2Real Transfer
The trained policies were first deployed and verified in simulation, where we observed interesting emergent behaviors among the agents. The agents strategically slowed down or steered away from each other to avoid collision (refer Fig. 4).
@@
Finally, it is worth mentioning that the closeness between sim and real metrics estimates the sim2real gap, which was least for LDR-MARL (4.12%) followed by NDR-MARL (9.38%), FGM (16.01%), and HDR-MARL (33.85%).
@@ IV-B2 Deployment and Sim2Real Transfer
The trained policies were first deployed and verified in simulation, where we observed interesting adversarial behaviors (e.g., blocking, baiting, overtaking, etc.). These behaviors conveyed that the agents were explicitly competing, while implicitly coordinating to avoid collisions (refer Fig. 7).
@@
Finally, the sim2real gap was least for LDR-MARL (2.88%) followed by NDR-MARL (6.88%), HDR-MARL (8.98%), and FGM (13.48%).
@@ V Conclusion
This work identified two pain points in training and deploying MARL systems, and attempted to address them by proposing a scalable and parallelizable digital twin framework. Two representative case studies were formulated to support the claims: a 4-agent collaborative intersection traversal problem and a 2-agent adversarial head-to-head racing problem. The two problems were deliberately formulated with distinct observation spaces and reward functions, but more importantly, also the learning architecture (vanilla MARL vs. demonstration-guided MARL). We analyzed the training metrics in each case and also noted the non-linear effect of agent/environment parallelization on the training time, with a hardware/software-specific point of diminishing return. Finally, we presented a mixed-reality sim2real transfer of the trained policies using a single physical vehicle, which was immersed within the proposed digital twin framework to interact with its virtual peers in a virtual environment.
Future avenues of research include analyzing the effect of different communication frameworks and protocols on digital twinning, formulation of physics-guided MARL problems, and scaling the deployments in terms of the number and size of the agents.